OpenStudy (anonymous):

the equation 5x + 2y=6 describes a line and is written in standard form (a) solve the variable y (b) use this result to find y with x = -8 i think i have this part right -5x + 3 2 the 5 and 2 is a fraction not sure how to use the draw :) just not sure where to go from here? thanks in advance

OpenStudy (anonymous):

that is in y=mx+b format

OpenStudy (anonymous):

be right back

OpenStudy (anonymous):

still new to this so not understanding your answer pmcafee

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

okay lets start from the start.............. what did you say y is

OpenStudy (anonymous):

2Y

OpenStudy (anonymous):

Your solution is good. -5/2x + 3. Now substitute -8 for x. y = -5/2(-8) + 3 y = ?

OpenStudy (anonymous):

have to simplify that

OpenStudy (anonymous):

23 thanks